# A geometric formulation to measure global and genuine entanglement in three-qubit systems

### v1

@@ 1 Introduction
Entanglement is the most interesting nonclassical correlation of multipartite quantum systems. It represents an important resource in quantum computing, quantum information processing and quantum teleportation [1, 2, 3]. However, the characterization and quantification of multipartite entanglement is still an open question [4, 5]. Even in the three-qubit system case it has been pointed out that different forms of entanglement may be present [6, 7, 8, 9]. This can be seen from the fact that there is no unified notion of a maximally entangled state. In fact, considering the most widely used measures [10], different requirements comprise different characteristics of the nonlocal properties that a given system must satisfy to exhibit maximal entanglement.
In this sense, entanglement measures based on the geometry of the state space or the appropriate projective space are of particular interest. For example, the geometric measure of entanglement (GME) evaluates the distance from a target state to its closest separable state in a given Hilbert space [11, 12]. Although its immediate geometric interpretation, the GME demands a non-trivial optimization so it becomes a nondeterministic polynomial (NP) problem [13]. Indeed, the amount of information that must be processed suggests that sooner or later it will be inevitable to resort to numerical methods [12]. Other examples includes the entanglement of minimum bipartite entropy [14], which quantifies the distance of a given state with its nearest state with no three-way entanglement, and robustness [15].
With respect to the geometry of projective spaces, for permutation invariant states, the Majorana representation leads to the identification of maximal symmetric $n$-qubit states [16, 17]; it has also been reported an entanglement measure associated with the barycenter of the Majorana constellation [16]. More recently, the triangle whose edges correspond to the squared bipartite concurrence of a three-qubit system has been considered as a useful tool to quantify entanglement [18]. It has been shown that the genuinely multipartite concurrence defined in [19] is exactly the square root of the shortest edge length of such a triangle, and that its perimeter is nothing else than the global entanglement measure considered in [20, 21]. Remarkably, the concurrence triangle area, computed through the Heron’s formula, is associated with a genuine entangled measure referred to as the concurrence fill [18].
@@
The most striking feature of our approach is to find that the projection and rejection of a given point of the polytope into the corresponding biseparable subsets yields a way to quantify entanglement. The former identifies global entanglement and the latter indicates genuine entanglement.
@@ 2 Qualitative characterization of entanglement
$|\psi\rangle=b_{0}|000\rangle+b_{1}e^{i\theta}|100\rangle+b_{2}|101\rangle+b_{3}|110\rangle+b_{4}|111\rangle,\quad b_{\ell}\geq 0,\quad\sum_{\ell=0}^{4}b_{\ell}^{2}=1,$ (1)
with $0\leq\theta\leq\pi$, we will pay attention to the vector $\vec{\lambda}_{\psi}=\left(\lambda_{1},\lambda_{2},\lambda_{3}\right)^{T}$, where $\lambda_{k}$ is the smallest eigenvalue of the density matrix $\rho_{k}$ associated to the $k$th qubit.
Looping through all allowed values of $b_{\ell}$ and $\theta$ in (1), vector $\vec{\lambda}_{\psi}$ localizes the points of a convex polytope $\mathcal{P}\subset\mathbb{R}^{3}$ that encodes the degree of entanglement between the different parts of a tripartite qubit system [24, 23, 25]. Therefore, we consider the mapping
@@
In general, depending on the number of parameters $b_{\ell}\neq 0$ that characterize the linear combination (1), there is a classification of entangled states [7, 8] that is summarized in Table 1. Next, based on the results reported in [24], we show that this classification results in the identification of concrete convex subsets of $\mathcal{P}$ under the mapping $\Lambda$.
@@
5 $\{|000\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle\}$ $\mathcal{P}$
Table 1: The number of basis elements of $\mathcal{H}=\mathcal{H}_{2}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{2}$ that are included in the linear superposition (1) gives rise to a classification of three-qubit states [7, 8]. Some of the representative states generated by such basis elements are associated with specific convex subspaces of the entanglement–polytope $\mathcal{P}\subset\mathbb{R}^{3}$ shown in Figure 1, see [24] and discussion in the main text. The class generated by any base element of $\mathcal{H}$ contains type 1 states; in the table we use $|000\rangle$ as a generic case.
The simplest class, called type 1, refers to superpositions (1) with only one coefficient different from zero. Any state of this type is fully separable, hereafter we write $|\psi_{\operatorname{sep}}\rangle=\ket{\phi_{1}}\ket{\phi_{2}}\ket{\phi_{3}}$, with $|\phi_{i}\rangle\in\mathcal{H}_{2}$. The latter implies that the smallest eigenvalue $\lambda_{k}$ of $\rho_{k}$ is equal to zero for all $i=1,2,3$. Thus, the entire set of vectors $|\psi_{\operatorname{sep}}\rangle$ (represented by $|000\rangle$ in Table 1) is associated with the vertex $\vec{S}\in\mathcal{P}$, which is a 0-dimensional convex set.
Type 2 includes linear superpositions (1) with two coefficients different from zero. We identify two different subclasses, called 2a and 2b.
@@
Type 2b includes superpositions (1) that are characterized by the pair $\{b_{0},b_{4}\}$. These states are associated with the one-dimensional convex subset generated by the extremal points $\vec{S}$ and $\vec{G}$ (the line-segment $\overline{SG}$). Explicitly, $\vec{\lambda}_{\psi_{2b}}=(1-2\lambda)\vec{S}+2\lambda\vec{G}$, where $\lambda=(1-\sqrt{1-4b_{0}^{2}b_{4}^{2}})/2$. The Greenberger–Horne–Zeilinger (GHZ) state
@@
States of type 3 involve three coefficients different from zero in the linear superposition (1). They are associated with two-dimensional convex subsets of $\mathcal{P}$, as we indicate below.
The type 3a refers to vectors that can be written in the form
@@
On the other hand, the type 3b includes three different families of vectors, given by the expressions
$\begin{array}[]{l}|\psi_{\rm 3b-1}\rangle=b_{0}|000\rangle+b_{1}e^{i\theta}|100\rangle+b_{4}|111\rangle,\\[6.88889pt] |\psi_{\rm 3b-2}\rangle=b_{0}|000\rangle+b_{2}|101\rangle+b_{4}|111\rangle,\\[6.88889pt] |\psi_{\rm 3b-3}\rangle=b_{0}|000\rangle+b_{3}|110\rangle+b_{4}|111\rangle.\end{array}$ (11)
@@
The type 4, referring to superpositions (1) that admit four non-zero coefficients, corresponds to three-dimensional convex subsets of $\mathcal{P}$ that are subdivided into four different clases.
@@
Type 4b is subdivided into two different categories, called b-1 (with $b_{2}=0$) and b-2 (with $b_{3}=0$). They are respectively mapped into a subset of $SB_{1}B_{3}G$ and $SB_{1}B_{2}G$ [25].
For states of type 4c (with $b_{1}=0$), let us consider the linear superposition
@@
The type 4d includes vectors that are expressed in an alternative canonical form [35],
This class of states is associated with points along the entire polytope $\mathcal{P}$. The relationship between types 4b, 4c and 4d, has been numerically studied in [25].
@@ 3 Quantitative characterization of entanglement
Figure 7: The arithmetic mean $\xi(\psi)$ introduced in Eq. (19) vanishes for fully separable states and is normalized to 1 for the GHZ–state. (a) The points $\vec{\lambda}_{\psi}\in\mathcal{P}$ average better values $\xi$ as they get closer to the vertex $\vec{G}\in\mathcal{P}$. (b) Condition $\xi(\psi)=\xi_{0}$, with $\xi_{0}=\operatorname{constant}$, defines planes transverse to the line segment $\overline{SG}$ whose points represent states $|\psi\rangle\in\mathcal{H}$ with the same degree of entanglement. The $\xi_{0}$–planes shown in the figure correspond to $\xi_{0}=\nicefrac{{1}}{{3}},\nicefrac{{2}}{{3}}$, and $\nicefrac{{5}}{{6}}$.
@@
Figure 8 shows the distribution of points $\vec{\lambda}_{\psi}$ according to the function $\mu$.
Figure 8: The transverse–distance measure $\mu(\psi)$ introduced in (21) gives 0 for separable states and 1 for the GHZ–state. In addition, this also gives 0 for bi-separable states, so it discriminates bi-separability from genuine entanglement. Notably, $\mu$ gives the same value as $\xi$ for the W–state, which confirms that $|W\rangle$ is $\xi=\mu=1/3$ distant from $|GHZ\rangle$.
@@
Figure 9 shows the values provided by the measures $\xi$ and $\mu$ for states $|\psi_{\operatorname{sep}}\rangle$, $|\psi_{\operatorname{bi}}\rangle=|0\rangle_{i}|\psi^{+}\rangle_{jk}$, $|W\rangle$ and $|GHZ\rangle$. These results are compared with the values that can be obtained in terms of the concurrence–fill $F_{123}$ [18] and the genuine multipartite concurrence $\mathcal{C}_{\operatorname{GME}}$ [19]. With exception of $\xi$, these measures give 0 for bi-separable states. That is, $\xi$ does not discriminate bi-separability from genuine entanglement, so it is a measure of global entanglement. On the other hand, among these measures, $\mu$ averages the lowest results for the four states. In other words, $\mu$ defines a lower bound for measuring genuine entanglement.

### v2

@@ 1 Introduction
Entanglement is the most interesting nonclassical correlation of multipartite quantum systems. It represents an important resource in quantum computing, quantum information processing and quantum teleportation [1, 2, 3]. However, the characterization and quantification of multipartite entanglement is still an open question [4, 5]. Even in the three-qubit system case it has been pointed out that different forms of entanglement may be present [6, 7, 8, 9]. This can be seen from the fact that there is no unified notion of a maximally entangled state for more than two-qubits [6]. In fact, considering the most widely used measures [10], different requirements comprise different characteristics of the nonlocal properties that a given system must satisfy to exhibit maximal entanglement.
In this sense, entanglement measures based on the geometry of the state space or the appropriate projective space are of particular interest. For example, the geometric measure of entanglement (GME) evaluates the distance from a target state to its closest separable state in a given Hilbert space [11, 12]. Although its immediate geometric interpretation, the GME demands a non-trivial optimization so it becomes a nondeterministic polynomial (NP) problem [13]. Indeed, the amount of information that must be processed suggests that sooner or later it will be inevitable to resort to numerical methods [12]. Other examples include the entanglement of minimum bipartite entropy [14], which quantifies the distance of a given state with its nearest state with no three-way entanglement, and robustness [15].
With respect to the geometry of projective spaces, for permutation invariant states, the Majorana representation leads to the identification of maximal symmetric $n$-qubit states [16, 17]; it has also been reported an entanglement measure associated with the barycenter of the Majorana constellation [16]. More recently, the triangle whose edges correspond to the squared bipartite concurrence of a three-qubit system has been considered as a useful tool to quantify entanglement [18]. It has been shown that the genuinely multipartite concurrence defined in [19] is exactly the square root of the shortest edge length of such a triangle, and that its perimeter is nothing else than the global entanglement measure considered in [20, 21]. Remarkably, the concurrence triangle area, computed through the Heron’s formula, is associated with a genuine entangled measure referred to as the concurrence fill [18]. Geometric simplices (tetrahedra) can be used to quantify GME [22], a generalization of the Peres-Horodeckis criterion applies also for multiqubit, continuous-variable, and hybrid systems [23], and even entropic measures derived from experimental correlations are useful to quantify tripartite entanglement [24].
@@
The most striking feature of our approach is to find that the projection of a given point of the polytope along the edges that represent biseparable states encodes a quantification of both, global and genuine entanglement.
@@ 2 Qualitative characterization of entanglement
$|\psi\rangle=b_{0}|000\rangle+b_{1}e^{i\omega}|100\rangle+b_{2}|101\rangle+b_{3}|110\rangle+b_{4}|111\rangle,\quad b_{\ell}\geq 0,\quad\sum_{\ell=0}^{4}b_{\ell}^{2}=1,$ (1)
with $0\leq\omega\leq\pi$, we will pay attention to the vector $\vec{\lambda}_{\psi}=\left(\lambda_{1},\lambda_{2},\lambda_{3}\right)^{T}$, where $\lambda_{k}$ is the smallest eigenvalue of the density matrix $\rho_{k}$ associated to the $k$th qubit.
Looping through all allowed values of $b_{\ell}$ and $\omega$ in (1), vector $\vec{\lambda}_{\psi}$ localizes the points of a convex polytope $\mathcal{P}\subset\mathbb{R}^{3}$ that encodes the degree of entanglement between the different parts of a tripartite qubit system [27, 26, 28]. Therefore, we consider the mapping
@@
System (3) characterizes $\mathcal{P}$ as the geometric scenario where the marginal problem defined by the pure state $|\psi\rangle$ finds solutions with physical meaning [26]. That is, among all possible states of one-qubit, $\lambda_{k}$ in (3) selects those that can be achieved as the reduced one-qubit density matrix $\rho_{k}=\operatorname{Tr}_{i,j}\left(|\psi\rangle\langle\psi|\right)$, with $k\neq i\neq j$. No state of one-qubit that is not linked to $|\psi\rangle$ in this way is included in $\mathcal{P}$. Indeed, (3) is a system of triangle inequalities that completely characterize the possible reduced one-qubit states of $|\psi\rangle$ [29].
@@
Considering the separability properties of $|\psi\rangle$, as it is written in (1), and remembering that local-unitary transformations preserve entanglement properties [36], one can identify different entangled pure states in terms of the coefficients $b_{\ell}\neq 0$, see details in Appendix A. Our classification coincides with the results reported in [7, 8], and is summarized in Table 1.
For completeness, in Table 1 we have included a set of states that is not obtained directly from (1). These states are classified as type 4d and are represented by the linear combination [37]:
Note that unlike $|\psi\rangle$, these additional states embrace only four basis vectors. Their usefulness will become clear in the following sections.
Next, based on the results reported in [27], we show that this classification results in the identification of concrete convex subsets of $\mathcal{P}$ under the mapping $\Lambda$.
@@
5 $\{|000\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle\}$ $\subset\mathcal{P}$
Table 1: The number of basis elements of $\mathcal{H}=\mathcal{H}_{2}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{2}$ that are included in the linear superposition (1) gives rise to a classification of three-qubit states [7, 8]. Some of the representative states generated by such basis elements are associated with specific convex subspaces of the entanglement–polytope $\mathcal{P}\subset\mathbb{R}^{3}$ shown in Figure 1, see [27] and discussion in the main text. The class of fully separable states is represented by $|000\rangle$ as a generic case, see details in Appendix A. States of type 4d are represented by the vector $|\psi_{4d}\rangle$ defined in (5).
The simplest class, called type 1, refers to fully separable states, hereafter written $|\psi_{\operatorname{sep}}\rangle=\ket{\phi_{1}}\ket{\phi_{2}}\ket{\phi_{3}}$, with $|\phi_{i}\rangle\in\mathcal{H}_{2}$. These states are generically represented by $|000\rangle$ in Table 1, since they are local-unitary equivalent to such state. Furthermore, in this case the smallest eigenvalue $\lambda_{k}$ of $\rho_{k}$ is equal to zero for all $i=1,2,3$. Thus, the entire set of vectors $|\psi_{\operatorname{sep}}\rangle$ is associated with the vertex $\vec{S}\in\mathcal{P}$, which is a 0-dimensional convex set.
Type 2 includes linear superpositions (1) with two coefficients different from zero as well as any other state $|\psi\rangle$ that can be transformed into any of these superpositions by means of local-unitary transformations, see Appendix A for details. We identify two different subclasses, called 2a and 2b.
@@
The generic type 2 state is characterized by the pair $\{b_{0},b_{4}\}$. This class of states is associated with the one-dimensional convex subset generated by the extremal points $\vec{S}$ and $\vec{G}$ (the line-segment $\overline{SG}$). Explicitly, $\vec{\lambda}_{\psi_{2b}}=(1-2\lambda)\vec{S}+2\lambda\vec{G}$, where $\lambda=(1-\sqrt{1-4b_{0}^{2}b_{4}^{2}})/2$. The Greenberger–Horne–Zeilinger (GHZ) state
@@
Generic states of type 3 involve three coefficients different from zero in the linear superposition (1). They are associated with two-dimensional convex subsets of $\mathcal{P}$, as we indicate below.
The type 3a refers to generic vectors that can be written in the form
@@
On the other hand, the type 3b includes three different families of generic vectors, given by the expressions
$\begin{array}[]{l}|\psi_{\rm 3b-1}\rangle=b_{0}|000\rangle+b_{1}e^{i\omega}|100\rangle+b_{4}|111\rangle,\\[6.88889pt] |\psi_{\rm 3b-2}\rangle=b_{0}|000\rangle+b_{2}|101\rangle+b_{4}|111\rangle,\\[6.88889pt] |\psi_{\rm 3b-3}\rangle=b_{0}|000\rangle+b_{3}|110\rangle+b_{4}|111\rangle.\end{array}$ (12)
@@
The generic type 4 states, referring to superpositions (1) that admit four non-zero coefficients, correspond to three-dimensional convex subsets of $\mathcal{P}$ that are subdivided into four different classes.
@@
Type 4b is subdivided into two different categories, called b-1 (with $b_{2}=0$) and b-2 (with $b_{3}=0$). They are respectively mapped into a subset of $SB_{1}B_{3}G$ and $SB_{1}B_{2}G$, see Figure 5.
(a) Type 4b-1
(b) Type 4b-2
Figure 5: Type 4b states. Following [28], the values of the Fourier coefficients have been discretized to identify the image of the states.
For states of type 4c (with $b_{1}=0$), the image covers region $SB_{1}B_{2}B_{3}\cup SB_{2}B_{3}G$. To identify states that are mapped to the faces of the upper tetrahedron of $\mathcal{P}$ we fix $b_{0}=1/\sqrt{2}$, and consider the linear superposition
@@
The generic type 4d state is expressed in an alternative canonical form [37], given in (5). This class of states is associated with points along the entire polytope $\mathcal{P}$. The relationship between types 4b, 4c, 4d and the polytope has been numerically studied in [28].
@@ 3 Quantitative characterization of entanglement
Global entanglement is usually quantified in terms of bipartite entanglement measures [15, 20, 36]. An example is the measure $Q$ introduced in [20], which is the sum of concurrences between a single qubit and the remaining qubits in the system. However, $Q$ does not distinguish between fully inseparable states and entangled states that are separable according to some set of subsystems [40]. On the other hand, one says that a pure state is genuinely entangled if it cannot be written as the product of simpler entangled states [19], the canonical examples are states $|GHZ\rangle$ and $|W\rangle$. In general, a given measure quantifies genuine entanglement if (a) it returns the result zero for fully separable states as well as for bi-separable states and (2) it is positive for non-biseparable states [19, 18, 22].
@@
Figure 8: The arithmetic mean $\xi(\psi)$ introduced in Eq. (19) vanishes for fully separable states and is normalized to 1 for the GHZ–state. (a) The points $\vec{\lambda}_{\psi}\in\mathcal{P}$ average higher values $\xi$ as they get closer to the vertex $\vec{G}\in\mathcal{P}$. (b) Condition $\xi(\psi)=\xi_{0}$, with $\xi_{0}=\operatorname{constant}$, defines planes transverse to the line segment $\overline{SG}$ whose points represent states $|\psi\rangle\in\mathcal{H}$ with the same degree of entanglement. The $\xi_{0}$–planes shown in the figure correspond to $\xi_{0}=\nicefrac{{1}}{{3}},\nicefrac{{2}}{{3}}$, and $\nicefrac{{5}}{{6}}$.
@@
Figure 9(a) shows the distribution of points $\vec{\lambda}_{\psi}$ according to the function $\mu$. Similar to the previous case, condition $\mu(\psi)=\mu_{0}=\operatorname{const}$ provides a collection of points that represent pure states $|\psi\rangle$ with the same degree of entanglement. Figure 9(b) shows the regions of $\mathcal{P}$ associated with two different values of $\mu_{0}$.
Figure 9: (a) The transverse–distance measure $\mu(\psi)$ introduced in (21) gives 0 for separable states and 1 for the GHZ–state. In addition, this also gives 0 for bi-separable states, so it discriminates bi-separability from genuine entanglement. Notably, $\mu$ gives the same value as $\xi$ for the W–state, which confirms that $|W\rangle$ is $\xi=\mu=1/3$ distant from $|GHZ\rangle$. (b) Regions of constant measure $\mu(\psi)=\mu_{0}$ for $\mu_{0}=4/5$ (closest to vertex $G$) and $\mu_{0}=1-\sqrt{13}/5$ (closest to vertex $S$).
@@
Figure 10 shows the values provided by the measures $\xi$ and $\mu$ for states $|\psi_{\operatorname{sep}}\rangle$, $|\psi_{\operatorname{bi}}\rangle=|0\rangle_{i}|\psi^{+}\rangle_{jk}$, $|W\rangle$ and $|GHZ\rangle$. These results are compared with the values that can be obtained in terms of the concurrence–fill $F_{123}$ and the genuine multipartite concurrence $\mathcal{C}_{\operatorname{GME}}$. The former measure is defined as [18]:
$F_{123}(\psi)=\left[8Q(\psi)\left(\frac{3}{2}Q(\psi)-\mathcal{C}_{1}^{2}\right)\left(\frac{3}{2}Q(\psi)-\mathcal{C}_{2}^{2}\right)\left(\frac{3}{2}Q(\psi)-\mathcal{C}_{3}^{2}\right)\right]^{1/4},$
where $Q\left(\psi\right)=\frac{1}{3}\sum_{i=1}^{3}\mathcal{C}_{i}^{2}\left(\psi\right)$. The $I$ concurrence $\mathcal{C}_{i}$ is a measure of entanglement between the $i$th qubit and the subsystem composed of qubits $j$ and $k$ [41]. In turn, the measure $\mathcal{C}_{GME}$ is defined as [19]:
$C_{GME}(\psi)=\min\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}.$
With exception of $\xi$, these measures give 0 for bi-separable states. That is, $\xi$ does not discriminate bi-separability from genuine entanglement, so it is a measure of global entanglement. On the other hand, among these measures, $\mu$ averages the lowest results for the four states. In other words, $\mu$ defines a lower bound for measuring genuine entanglement.
@@ 4 The inverse problem
We can also pose the inverse problem in another context. Suppose you are interested in determining a collection of points that describe a given path in $\mathcal{P}$. Say, the path follows one of the surfaces determined by $\mu=\mu_{0}=\operatorname{const}$. It can be shown that a particular solution is obtained after making $b_{\ell}=1/\sqrt{5}$, $\ell=0,1,2,3,4$, in (1). Looping through all the domain $[0,\pi]$ of $\omega$, we obtain the path shown in Figure 13(a). In a given interval of $\omega$, the path overflows the $\mu_{0}$-region towards vertex $S$. That is, entanglement is constant for $\omega\in[\omega_{0},\pi]$ and decreases as $\omega$ goes from $\omega_{0}$ to zero. Our assertion is verified in Figure 13(b), where we show the measures $\mu$, $\xi$, $F_{123}$ and $\mathcal{C}_{GME}$ for the $\omega$-dependent solution we have found to the above inverse problem.
Figure 13: A particular solution to the inverse problem of finding a set of points in $\mathcal{P}$ that follows a path with entanglement measure $\mu=\mu_{0}=\operatorname{const}$, see Figure 9(b), is obtained after making $b_{\ell}=1/\sqrt{5}$ in (1). The pure state $|\psi\rangle$ is parameterized by the phase $0\leq\omega\leq\pi$ and is projected onto the path $\vec{\lambda}_{\psi(\omega)}$ in $\mathcal{P}$ (a) $\vec{\lambda}_{\psi(\omega)}$ starts and finishes at $\left(\nicefrac{{\mu_{0}}}{{2}},\varepsilon,\varepsilon\right)$ and $\left(\nicefrac{{\mu_{0}}}{{2}},\nicefrac{{2}}{{5}},\nicefrac{{2}}{{5}}\right)$, respectively. Here $\mu_{0}=1-\nicefrac{{\sqrt{13}}}{{5}}$ and $\varepsilon=\tfrac{1}{2}\left(1-\nicefrac{{\sqrt{17}}}{{5}}\right)$. In the interval $[0,\omega_{0}]$, with $\omega_{0}=\pi/3$, the path overflows the $\mu_{0}$-region. (b) Measures $\mu$, $\xi$, $F_{123}$ and $\mathcal{C}_{GME}$ as functions of $\omega$. Only $\mu$ and $\mathcal{C}_{GME}$ maintain constant in the interval $[\omega_{0},\pi]$, with values $\mu=1-\nicefrac{{\sqrt{13}}}{{5}}$ and $\mathcal{C}_{GME}=\nicefrac{{12}}{{25}}$.